OpenStudy (anonymous):

Solve the equation: log 2(x-3)=5 A. 32 B.5 c.No solution D.35

OpenStudy (anonymous):

2^5=x-3 32=x-3 35=x Answer is D.

OpenStudy (anonymous):

Thanks,

OpenStudy (anonymous):

welcome